# Comprehensive Evaluation of Prototype Neural Networks

### v2

@@ 7 Results
The used networks are ProtoPNet [16], ProtoPNet Pruned [16], ProtoPool [20], and PIPNet [21]. The pruned version of ProtoPNet was created using the pruning strategy stated in the original paper.
@@ 7.1 General and Compactness
PIPNet achieves the best overall results in both predictive performance and compactness. This illustrates the potential of contrastive learning for prototype-based methods. The outcome is intuitive, as contrastive learning naturally aligns with the goal of learning high-level semantic prototype representation. Furthermore, the hard constraint imposed on the classification layer does not negatively affect performance, indicating that PIPNet does not require a separate fine-tuning phase. In contrast, other models that rely on fine-tuning often learn additional, more complex, relationships to increase predictive performance in this phase. PIPNet is also the only model capable of reducing Global Size during training without the need for external pruning strategies. This highlights the effectiveness of the softmax-based constraint in limiting the number of active prototypes, even while the tanh loss encourages all prototypes to be active at some point during training. Moreover, the consistently small Local Size relative to Global Size indicates that this design robustly limits prototype usage per sample, thereby enhancing interpretability without sacrificing accuracy.
@@ 7.2 Contrastivity
All contrastivity results in Figure 3 were normalized using either fixed metric boundaries or the maximum values observed across models. The metric axes, where lower values indicate better performance, are inverted to consistently increase the coverage area of the more powerful models.
@@ 7.3 Continuity
The design of ProtoPool also does not include such a mechanism. However, it appears learning shared prototypes promote the extraction of more robust features under the right circumstances. This is reflected in the superior performance on the NICO and AWA2 dataset compared to ProtoPNet. A noteworthy finding is the high PLC score relative to other similarity-map metrics, such as PALC. This suggests that while the location of maximum activation may vary, the broader activation patterns remain relatively stable. Furthermore, ProtoPool performs particularly well on the AWA2 dataset, indicating that the multi-label classification loss contributes to learning more robust shared prototypes. Another key observation is that better PRC scores do not necessarily translate into improved CRC scores. This again highlights the complex and sensitive nature of the classification process from the trained ProtoPool models, which is in agreement with the findings from the compactness evaluation.
@@ 7.4 Covariate Complexity
The ProtoPool model exhibits the highest Background Overlap, indicating that it is particularly affected by this issue. This observation aligns with its overall performance and compactness results on the CUB200 dataset, which suggest that the learned prototypes lack sufficient class-discriminative information and are of comparatively low quality.
@@ 8 Discussion
On CUB200 and Cars196, our accuracies are 5–10% below the original reports, most likely due to halving prototype channels (256→128) for more stability under our training setup. In addition, only the classification loss was class-weighted to counter imbalance; cluster/separation losses were not. Pruning produced only marginal gains in interpretability, falling short of the clearer pruning effects reported by Chen et al. [16]. Compared to Wang et al. [64], our PRP-based visualizations yield smaller, sparser activations than the original upsampling method, explaining their higher Object Overlap measures.
@@
Similar to ProtoPNet, our ProtoPool scores on CUB200 and Cars196 are 5–10% below the original, that we explain by the same channel reduction and class-imbalance handling.
@@
However, visual explanations (Figures 6 and 7) reveal issues similar to those observed in ProtoPNet and ProtoPool. While prototype concepts can often be guessed in single-image visualizations, the nearest feature comparisons again reveal mismatches. Interestingly, PIPNet focuses more on background regions in the NICO and AWA2 datasets, showing the opposite trend of ProtoPool, despite both models employing shared prototypes. Relative to Nauta et al. [21], our PIPNet models underperform by 8%, which we primarily attribute to the lower feature map resolution used here ($7{\times}7$ vs. $28{\times}28$) to ensure a uniform comparison across methods. Despite this gap, our compactness results mirror the original findings: comparable reductions in global size and high sparsity in the classification layer. PRP visualizations align best with PIPNet, though object-part consistency drops notably compared to the purity metric from the original study that focuses only on the top-10 nearest training images. On AWA2, PIPNet attains the highest accuracy. However, performance is likely suppressed by highly imbalanced labels and a multi-label margin loss that cannot incorporate class weights.
Overall, our evaluation focused on diverse model aspects relevant to interpretability, with mixed results across models and reappearing patterns in different evaluation techniques from individual architectures.
@@ 9 Conclusion
We examined the evaluation of prototype-based neural networks, drawing on the foundational research by Nauta and Seifert [19] and Nauta et al. [67]. The Co-12 properties identified by Nauta et al. [67] provide a robust framework for assessing the quality of explanations in XAI methods. We complement these using several new metrics and provide a comprehensive evaluation using four datasets. Our python library integrates the existing evaluation techniques in the Co-12 framework and offers a practical toolkit for validating, benchmarking, and comparing part-based prototype networks.

### v3

@@ 7 Results
The used networks are ProtoPNet Chen et al. [2019], ProtoPNet Pruned Chen et al. [2019], ProtoPool Rymarczyk et al. [2022], and PIPNet Nauta et al. [2023a]. The pruned version of ProtoPNet was created using the pruning strategy stated in the original paper.
@@ 7.1 General and Compactness
PIPNet achieves the best overall results in both predictive performance and compactness. This illustrates the potential of contrastive learning for prototype-based methods. The outcome is intuitive, as contrastive learning naturally aligns with the goal of learning a high-level semantic prototype representation. Furthermore, the hard constraint imposed on the classification layer does not negatively affect performance, indicating that PIPNet does not require a separate fine-tuning phase. In contrast, other models that rely on fine-tuning often learn additional, more complex, relationships to increase predictive performance in this phase. PIPNet is also the only model capable of reducing Global Size during training without the need for external pruning strategies. This highlights the effectiveness of the softmax-based constraint in limiting the number of active prototypes, even while the tanh loss encourages all prototypes to be active at some point during training. Moreover, the consistently small Local Size relative to Global Size indicates that this design robustly limits prototype usage per sample, thereby enhancing interpretability without sacrificing accuracy.
@@ 7.2 Contrastivity
All contrastivity results in Figure 3 were normalized using either fixed metric boundaries or the maximum values observed across models. The metric axes, where lower values indicate better performance, are inverted to consistently increase the coverage area of the better performing models.
@@ 7.3 Continuity
The design of ProtoPool also does not include such a mechanism. However, it appears learning shared prototypes promote the extraction of more robust features under the right circumstances. This is reflected in the superior performance on the NICO and AWA2 dataset compared to ProtoPNet. A noteworthy finding is the high PLC score relative to other similarity-map metrics, such as PALC. This suggests that while the location of maximum activation may vary, overall the broader activation patterns remain relatively stable. Furthermore, ProtoPool performs particularly well on the AWA2 dataset, indicating that the multi-label classification loss contributes to learning more robust shared prototypes. Another key observation is that better PRC scores do not necessarily translate into improved CRC scores. This again highlights the complex and sensitive nature of the classification process from the trained ProtoPool models, which is in agreement with the findings from the compactness evaluation.
@@ 7.4 Covariate Complexity
The ProtoPool model exhibits the highest Background Overlap, indicating that it is particularly affected by this issue. This observation aligns with the model’s overall performance and compactness results on the CUB200 dataset, which suggest that the learned prototypes lack sufficient class-discriminative information and are of comparatively low quality.
@@ 8 Discussion
On CUB200 and Cars196, our accuracies are 5–10% below the original reports, most likely due to halving prototype channels (256→128) for more stability under our training setup. In addition, only the classification loss was class-weighted to counter imbalance; cluster/separation losses were not. Pruning produced only marginal gains in interpretability, falling short of the clearer pruning effects reported by Chen et al. Chen et al. [2019]. Compared to Wang et al. Wang et al. [2023], our PRP-based visualizations yield smaller, sparser activations than the original upsampling method, explaining their higher Object Overlap measures.
@@
Similar to ProtoPNet, our obtained ProtoPool scores on CUB200 and Cars196 are 5–10% below the original, that we explain by the same channel reduction and class-imbalance handling.
@@
However, visual explanations (Figures 6 and 7) reveal issues similar to those observed in ProtoPNet and ProtoPool. While prototype concepts can often be guessed in single-image visualizations, the nearest feature comparisons again reveal mismatches. Interestingly, PIPNet focuses more on background regions in the NICO and AWA2 datasets, showing the opposite trend of ProtoPool, despite both models employing shared prototypes. Relative to Nauta et al. Nauta et al. [2023a], our PIPNet models underperform by 8%, which we primarily attribute to the lower feature map resolution used here ($7{\times}7$ vs. $28{\times}28$) to ensure a uniform comparison across methods. Despite this gap, our compactness results mirror the original findings: comparable reductions in global size and high sparsity in the classification layer. PRP visualizations align best with PIPNet, though object-part consistency drops notably compared to the purity metric from the original study that focuses only on the top-10 nearest training images. On AWA2, PIPNet attains the highest accuracy. However, performance is likely suppressed by highly imbalanced labels and a multi-label margin loss that cannot incorporate class weights.
Overall, our evaluation focused on diverse model aspects relevant to interpretability, with mixed results across models and reappearing patterns in different evaluation techniques from individual architectures. We see that models generally struggle to learn semantically distinct prototypes and are more focused on lower level image characteristics and shapes. This often results in a cramped latent space and similar activation locations, facilitating a complex relation between prototypes and classes. It is evident that learning semantic prototypes is a complex task, that current architectures struggle to consistently achieve. Other approaches that promote some of the evaluated properties in a more direct matter like contrastive learning approaches seem to be able to mitigate this problem well, as seen by the PIPNet model.
@@ 9 Conclusion
We examined the evaluation of prototype-based neural networks, drawing on the foundational research by Nauta and Seifert Nauta and Seifert [2023] and Nauta et al. Nauta et al. [2023b]. The Co-12 properties identified by Nauta et al. Nauta et al. [2023b] provide a robust framework for assessing the quality of explanations in XAI methods. We complement these using several new metrics and provide a comprehensive evaluation using four datasets. Our python library integrates the existing evaluation techniques in the Co-12 framework and offers a practical toolkit for validating, benchmarking, and comparing part-based prototype networks.
@@ Acknowledgments
Declarations
Competing Interests
The authors declare that they have no competing interests related to the subject matter discussed in this manuscript.
Data Availability
The datasets generated and analyzed during the current study are publicly available. Further information, including access links, can be found in the repository at https://github.com/uos-sis/quanproto.